# Discrete conformal structures on surfaces with boundary (II)—Rigidity and Existence

### v1

@@ 3.1. Admissible space of the discrete conformal structure (A1)
If $\alpha_{r}=1,\alpha_{s}=-1$, then $\cosh l_{rs}=-\sqrt{(1+e^{2f_{r}})(1-e^{2f_{s}})}+\eta_{rs}e^{f_{r}+f_{s}}=-\coth u_{r}\tanh u_{s}+\eta_{rs}\frac{-1}{\sinh u_{r}}\frac{1}{\cosh u_{s}}>1$. Since $u_{r},u_{s}\in\mathbb{R}_{<0}$, then $\sinh(u_{r}+u_{s})>-\eta_{rs}$ and hence $u_{r}+u_{s}>\mathrm{arcsinh}\ (-\eta_{rs})$.
In summary, the admissible space $\mathcal{U}_{ijk}(\eta)$ of $u$ can be rewritten as
(35) $\mathcal{U}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{N_{1}}\times\mathbb{R}_{<0}^{3-N_{1}}|u_{r}+u_{s}>C(\eta_{rs}),\ \forall\{r,s\}\subseteq\{i,j,k\}\},$
where $C(\eta_{rs})$ is a constant that depends on $\eta_{rs}$.
The above arguments can imply the following theorem.
@@
Suppose $(\Sigma,\mathcal{T},\alpha,\eta)$ is a weighted triangulated surface with boundary, and the weights $\alpha:B\rightarrow\{-1,0,1\}$ and $\eta\in\mathbb{R}_{>0}^{E}$ satisfying $\eta_{ij}>\alpha_{i}\alpha_{j}$ if $\alpha_{i}=\alpha_{j}$ for two adjacent boundary components $i,j\in B$. Then the admissible space $\mathcal{U}_{ijk}(\eta)$ is a convex polytope. As a result, the admissible space
$\mathcal{U}(\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\eta)$
is a convex polytope.
3.2. Negative definiteness of Jacobian
For simplicity, we set $l_{r}=l_{st},\ \theta_{r}=\theta^{st}_{r}$ and $A=\sinh l_{r}\sinh l_{s}\sinh\theta_{t}$, where $\{r,s,t\}=\{i,j,k\}$. Define
@@ 3.2. Negative definiteness of Jacobian
Then $\det Q_{1}=-1+2\cosh\theta_{i}\cosh\theta_{j}\cosh\theta_{k}+\cosh^{2}\theta_{i}+\cosh^{2}\theta_{j}+\cosh^{2}\theta_{k}>0$. By calculations similar to (16), we have
@@
For any $\alpha_{i}\in\{-1,0,1\}$, set $S_{i}=e^{f_{i}}$, $C_{i}=\sqrt{1+\alpha_{i}e^{2f_{i}}}$, then $S_{i},C_{i}>0$ and $C_{i}^{2}-\alpha_{i}S_{i}^{2}=1$. By (34), we have
@@
Under the same assumptions as in Theorem 3.1. For a right-angled hyperbolic hexagon $\{ijk\}\in F$ on the admissible space $\mathcal{U}_{ijk}(\eta)$, the Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is non-degenerate, symmetric and negative definite.
@@
Since $d>0$ for the discrete conformal structure (A1) in Theorem 1.6, then $\det Q_{2}>0$ by (39). Hence
@@
This implies the matrix $\Lambda_{ijk}$ is non-degenerate. By Lemma 2.22, the matrix $\Lambda_{ijk}$ is symmetric. Set $a_{rrs}:=C_{r}\cdot\coth d_{rs}>0$. Define
@@
To prove the negative definiteness of $\Lambda_{ijk}$, it suffices to prove $Q$ is positive definite. The formula (41) shows that $\det Q>0$. It remains to prove that the determinants of the $1\times 1$ and $2\times 2$ order principal minor of $Q$ are positive. Let $m_{rs}$ be the entry of $Q$ at $r$-th row and $s$-th column. Then
@@
Thus the determinants of the $1\times 1$ principal submatrices of $Q$ is $m_{11}>0$ by $a_{iik}>0,a_{iij}>0$. And the determinants of the $2\times 2$ principal submatrices of $Q$ is
@@
As a result, we have the following corollary.
@@
Under the same assumptions as in Theorem 3.1. The Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on the admissible space $\mathcal{U}(\eta)$.
@@ 3.3. Rigidity of the discrete conformal structure (A1)
The following theorem gives the rigidity of discrete conformal structure (A1), which is the case $\mathrm{(i)}$ in Theorem 1.9.
@@
Under the same assumptions as in Theorem 3.1. The function $f$ is determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$. In particular, the map from $f$ to $K$ is a smooth embedding.
@@
Theorem 3.1 and Theorem 3.2 imply that the following energy function
@@
for a right-angled hyperbolic hexagon $\{ijk\}\in F$ is a well-defined smooth function on $\mathcal{U}_{ijk}(\eta)$. Furthermore, $\mathcal{F}_{ijk}(u_{i},u_{j},u_{k})$ is a strictly concave function on $\mathcal{U}_{ijk}(\eta)$ with $\nabla_{u_{i}}\mathcal{F}_{ijk}=\theta^{jk}_{i}$. Define the function ~$\widetilde{\mathcal{F}}:\mathcal{U}(\eta)\rightarrow\mathbb{R}$ by
@@
By Corollary 3.3, ~$\widetilde{\mathcal{F}}(u)$ is a strictly concave function on $\mathcal{U}(\eta)$ with $\nabla_{u_{i}}\widetilde{\mathcal{F}}(u)=K_{i}$. According to Lemma 1.11, the map ~$\nabla\widetilde{\mathcal{F}}:\mathcal{U}(\eta)\rightarrow\mathbb{R}^{N}$ is a smooth embedding. Therefore, the map from $\{f\in\Omega(\eta)=\bigcap_{\{ijk\}\in F}\Omega_{ijk}(\eta)\}$ to $\{K\in\mathbb{R}^{N}_{>0}\}$ is a smooth injective map. Q.E.D.
3.4. Existence of discrete conformal structure (A1)
The following theorem characterizes the image of $K$.
@@ 3.4. Existence of discrete conformal structure (A1)
Under the same assumptions as in Theorem 3.1. Suppose $\alpha:B\rightarrow\{0,1\}$. Then the image of $K$ is >$\mathbb{R}^{N}_{>0}$.
@@
Under the same assumptions as in Theorem 3.1. For a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\Omega_{ijk}(\eta)$, if one of the following conditions is satisfied
@@
Then $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=0$, where $a,\ b,\ c$ are constants.
@@
For case $\mathrm{(i)}$, $\lim f_{i}=+\infty,\ \lim f_{j}=+\infty,\ \lim f_{k}=+\infty$. By (5), we have
@@
then $\cosh\theta^{jk}_{i}\rightarrow 1$ and hence $\theta_{i}^{jk}\rightarrow 0$. The same proofs are suitable for the cases $\mathrm{(ii)}$ and $\mathrm{(iii)}$, and we omit them here. Q.E.D.
@@
Under the same assumptions as in Theorem 3.1. For a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\Omega_{ijk}(\eta)$, if $\lim f_{i}=-\infty$, then $\lim f_{j}=\lim f_{k}=+\infty$. Furthermore, $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=+\infty$.
@@
According to the definition of the admissible space $\mathcal{U}_{ijk}(\eta)$ by (35), there is no subset $\{r,s\}\subseteq\{i,j,k\}$ such that $\lim u_{r}=-\infty$ and $\lim u_{s}=-\infty$ or $\lim u_{s}=c$, where $c$ is a constant. Since $\frac{\partial f_{r}}{\partial u_{r}}=\sqrt{1+\alpha_{r}e^{2f_{r}}}$ for $r\in\{i,j,k\}$, by Remark 2.23, if $\lim f_{i}=-\infty$, then $\lim u_{i}=-\infty$; if $\lim f_{j}=+\infty$, then $\lim u_{j}=+\infty$ with $\alpha_{j}=0$ and $\lim u_{j}=0^{-}$ with $\alpha_{j}=1$. Therefore, if $\lim f_{i}=-\infty$, then $\lim u_{i}=-\infty$ and hence $\lim u_{j}=\lim u_{k}=+\infty$. This implies $\alpha_{j}=\alpha_{k}=0$ and $\lim f_{j}=\lim f_{k}=+\infty$.
If $\alpha_{i}=0$, then $\cosh l_{ij}=-1+\eta_{ij}e^{u_{i}+u_{j}}$$\cosh l_{jk}=-1+\eta_{jk}e^{u_{j}+u_{k}}$ and $\cosh l_{ik}=-1+\eta_{ik}e^{u_{i}+u_{k}}$. Thus
@@
And $L_{2}>0$. Then $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, which implies $\theta^{jk}_{i}\rightarrow+\infty$.
@@
Similarly, $\cosh l_{ik}\rightarrow-1+2\eta_{ik}e^{u_{i}+u_{k}}$ and $\cosh l_{jk}=-1+\eta_{jk}e^{u_{j}+u_{k}}$. Similar to the proof of the case $\alpha_{i}=0$, we have $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, then $\theta^{jk}_{i}\rightarrow+\infty$.
@@
Similarly, $\cosh l_{ik}\rightarrow-1+2\eta_{ik}e^{u_{i}+u_{k}}$ and $\cosh l_{jk}=-1+\eta_{jk}e^{u_{j}+u_{k}}$. Similar to the proof of the case $\alpha_{i}=0$, we have $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, then $\theta^{jk}_{i}\rightarrow+\infty$. Q.E.D.
Proof of Theorem 3.5: Let the image of $K$ be $X$. To prove $X=\mathbb{R}^{N}_{>0}$, we will show that $X$ is both open and closed in >$\mathbb{R}^{N}_{>0}$. By the definition of curvature map $K$, we have $X\subseteq\mathbb{R}^{N}_{>0}$. Since the map from $f\in\Omega(\eta)$ to >$\mathbb{R}^{N}_{>0}$ is a smooth injective map, then $X$ is open in >$\mathbb{R}^{N}_{>0}$. It remains to prove that $X$ is closed in >$\mathbb{R}^{N}_{>0}$. We show that if a sequence $f^{(m)}$ in $\Omega(\eta)$ satisfies $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$, then there exists a subsequence, say $f^{(m)}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}$ is in $\Omega(\eta)$.
Suppose otherwise, there exists a subsequence $f^{(m)}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}$ is in the boundary of $\Omega(\eta)$. By (33), the boundary of the admissible space $\Omega(\eta)$ in $[-\infty,+\infty]^{N}$ consists of the following three parts
@@
There exists an ideal edge $\{rs\}\in E$ such that $l_{rs}=0$, i.e., $u_{r}+u_{s}=-M(\eta_{rs})$ for the admissible space $\mathcal{U}_{ijk}(\eta)$ defined by (35).
@@
For $\alpha\in\{0,1\}$, there exists $i\in B$ satisfying $f_{i}^{(m)}\rightarrow+\infty$.
@@
For $\alpha\in\{0,1\}$, there exists $i\in B$ satisfying $f_{i}^{(m)}\rightarrow-\infty$.
For case $\mathrm{(i)}$, in a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths $l_{ij},l_{jk},l_{ik}$ and opposite lengths of boundaries $\theta^{ij}_{k},\theta^{jk}_{i},\theta^{ik}_{j}$. Without loss of generality, we assume $l_{ij}\rightarrow 0$. By the cosine law,
@@
which contradicts the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$.
The case $\mathrm{(ii)}$ follows from Lemma 3.6. In fact, each generalized angle incident to the boundary component $i\in B$ converges to 0. Hence $\lim_{m\rightarrow+\infty}K^{(m)}=0$. This contradicts with the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$.
The case $\mathrm{(iii)}$ follows from Lemma 3.7. In fact, each generalized angle incident to the boundary component $i\in B$ converges to +$\infty$+ ∞. Hence $\lim_{m\rightarrow+\infty}K^{(m)}=+\infty$. This contradicts with the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$. Q.E.D.
@@
Note that Lemma 3.6 and Lemma 3.7 hold for $\alpha:B\rightarrow\{-1,0,1\}$. But Theorem 3.5 holds for $\alpha:B\rightarrow\{0,1\}$. If $\alpha_{i}=-1$, then the boundary of the admissible space $\Omega(\eta)$ in $[-\infty,+\infty]^{N}$ will include the case that there exists $i\in B$ satisfying $f_{i}^{(m)}\rightarrow 0^{-}$. We are unable to prove the closeness of $X$ in this case.

### v3

@@ 3.1. Admissible space of the discrete conformal structure (A1)
If $\alpha_{r}=1$ and $\alpha_{s}=-1$, then $\cosh l_{rs}=-\sqrt{(1+e^{2f_{r}})(1-e^{2f_{s}})}+\eta_{rs}e^{f_{r}+f_{s}}=-\coth u_{r}\tanh u_{s}+\eta_{rs}\frac{-1}{\sinh u_{r}}\frac{1}{\cosh u_{s}}>1$. Since $u_{r}\in\mathbb{R}_{<0}$ and $u_{s}\in\mathbb{R}_{<0}$, it follows that $\sinh\,(u_{r}+u_{s})>-\eta_{rs}$, and hence $u_{r}+u_{s}>\mathrm{arcsinh}\,(-\eta_{rs})$.
In summary, the admissible space $\mathcal{U}_{ijk}(\eta)$ can be rewritten as
(35) $\mathcal{U}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{N_{1}}\times\mathbb{R}_{<0}^{3-N_{1}}\mid u_{r}+u_{s}>C(\eta_{rs}),\ \forall\{r,s\}\subseteq\{i,j,k\}\},$
where $C(\eta_{rs})$ is a constant depending on $\eta_{rs}$.
@@
Suppose $(\Sigma,\mathcal{T},\alpha,\eta)$ is a weighted triangulated surface with boundary, where the weighs are given by $\alpha:B\rightarrow\{-1,0,1\}$ and $\eta\in\mathbb{R}_{>0}^{E}$ satisfying $\eta_{ij}>\alpha_{i}\alpha_{j}$ for any two adjacent boundary components $i,j\in B$. Then the admissible space $\mathcal{U}_{ijk}(\eta)$ is a convex polytope. As a result, the admissible space $\mathcal{U}(\eta)=\bigcap_{\{ijk\}\in F}\mathcal{U}_{ijk}(\eta)$ is also a convex polytope.
3.2. Negative definiteness of the Jacobian
For a hyper-ideal hyperbolic triangle $\{ijk\}\in F$, we introduce some simplifying notations: $l_{r}=l_{st}$, $\theta_{r}=\theta^{st}_{r}$, and $A=\sinh l_{r}\sinh l_{s}\sinh\theta_{t}$, where $\{r,s,t\}=\{i,j,k\}$. Define
@@ 3.2. Negative definiteness of the Jacobian
Then $\det Q_{1}=-1+2\cosh\theta_{i}\cosh\theta_{j}\cosh\theta_{k}+\cosh^{2}\theta_{i}+\cosh^{2}\theta_{j}+\cosh^{2}\theta_{k}>0$. Following calculations analogous to those in (15), we obtain
@@
For any $\alpha_{i}\in\{-1,0,1\}$, set $S_{i}=e^{f_{i}}$ and $C_{i}=\sqrt{1+\alpha_{i}e^{2f_{i}}}$, so $S_{i}>0$, $C_{i}>0$, and $C_{i}^{2}-\alpha_{i}S_{i}^{2}=1$. It follows from (34) that
@@
Under the same assumptions as those in Theorem 3.1, for a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\mathcal{U}_{ijk}(\eta)$, the Jacobian $\Lambda_{ijk}=\frac{\partial(\theta^{jk}_{i},\theta^{ik}_{j},\theta^{ij}_{k})}{\partial(u_{i},u_{j},u_{k})}$ is symmetric and negative definite.
@@
Since $d>0$ for the discrete conformal structure (A1) in Theorem 1.5, then $\det Q_{2}>0$ by (39). Consequently,
@@
This implies the matrix $\Lambda_{ijk}$ is non-degenerate. The symmetry of $\Lambda_{ijk}$ follows from Lemma 2.21. Set $a_{rrs}:=C_{r}\cdot\coth d_{rs}>0$. Define
@@
To prove the negative definiteness of $\Lambda_{ijk}$, it suffices to show that $Q$ is positive definite. Formula (41) shows that $\det Q>0$. It remains to verify that the $1\times 1$ and $2\times 2$ leading principal minors of $Q$ are positive. Let $m_{rs}$ denote the entry of $Q$ at $r$-th row and $s$-th column. Then
@@
Thus the $1\times 1$ leading principal minor of $Q$ is $m_{11}>0$ due to $a_{iik}>0$ and $a_{iij}>0$. And the $2\times 2$ leading principal minor of $Q$ is
@@
As a consequence, we obtain the following corollary.
@@
Under the same assumptions as those in Theorem 3.1, the Jacobian $\Lambda=\frac{\partial(K_{i},...,K_{N})}{\partial(u_{i},...,u_{N})}$ is symmetric and negative definite on $\mathcal{U}(\eta)$.
Proof.
The conclusion follows directly from Theorem 3.2 and the identity $\Lambda=\sum_{\{ijk\}\in F}\Lambda_{ijk}$. Specifically, each $\Lambda_{ijk}$ is extended to an $N\times N$ matrix via zero-padding, such that it acts non-trivially on a vector $(1,...,N)$ exclusively through the coordinates corresponding to the boundary components $i,j,k$ of the right-angled hyperbolic hexagon $\{ijk\}\in F$. Q.E.D.
@@ 3.3. Rigidity of the discrete conformal structure (A1)
The following theorem establishes the rigidity of the discrete conformal structure (A1), which corresponds to the rigidity part of Theorem 1.8 (i).
@@
Under the same assumptions as those in Theorem 3.1, the discrete conformal factor $f$ is uniquely determined by its generalized combinatorial curvature $K\in\mathbb{R}^{N}_{>0}$.
@@
Theorem 3.1 and Theorem 3.2 imply that for a right-angled hyperbolic hexagon $\{ijk\}\in F$, the following energy function
@@
is a well-defined smooth function on $\mathcal{U}_{ijk}(\eta)$. Furthermore, $\mathcal{F}_{ijk}(u_{i},u_{j},u_{k})$ is strictly concave on $\mathcal{U}_{ijk}(\eta)$ with $\nabla_{u_{i}}\mathcal{F}_{ijk}=\theta^{jk}_{i}$. Define the function ~$\widetilde{\mathcal{F}}:\mathcal{U}(\eta)\rightarrow\mathbb{R}$ by
@@
By Corollary 3.3, ~$\widetilde{\mathcal{F}}(u)$ is strictly concave on $\mathcal{U}(\eta)$ with $\nabla_{u_{i}}\widetilde{\mathcal{F}}(u)=K_{i}$. By Lemma 1.10, the gradient map ~$\nabla\widetilde{\mathcal{F}}:\mathcal{U}(\eta)\rightarrow\mathbb{R}^{N}$ is a smooth embedding. Note that $\mathcal{U}(\eta)$ and $\Omega(\eta)=\bigcap_{\{ijk\}\in F}\Omega_{ijk}(\eta)$ are diffeomorphic. Consequently, the map from $\{f\in\Omega(\eta)\}$ to $\{K\in\mathbb{R}^{N}_{>0}\}$ is a smooth injective map. Q.E.D.
3.4. Existence of the discrete conformal structure (A1)
The following theorem characterizes the image of $K$ for the discrete conformal structure (A1), which corresponds to the existence part of Theorem 1.8 (i).
@@ 3.4. Existence of the discrete conformal structure (A1)
Suppose $(\Sigma,\mathcal{T},\alpha,\eta)$ is a weighted triangulated surface with boundary, where the weights are given by $\alpha:B\rightarrow\{0,1\}$ and $\eta\in\mathbb{R}_{>0}^{E}$ satisfying $\eta_{ij}>\alpha_{i}\alpha_{j}$ for any two adjacent boundary components $i,j\in B$. The image of $K$ is >$\mathbb{R}^{N}_{>0}$.
@@
Under the same assumptions as those in Theorem 3.1, for a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\Omega_{ijk}(\eta)$, if one of the following conditions is satisfied
@@
where $a,b,c$ are constants. Then $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=0$.
@@
For the case (i), it follows from (5) that
@@
it follows that $\cosh\theta^{jk}_{i}\rightarrow 1$, which implies $\theta_{i}^{jk}\rightarrow 0$. The same arguments apply for the cases (ii) and (iii), the details of which are omitted here for brevity. Q.E.D.
@@
Under the same assumptions as those in Theorem 3.1, for a right-angled hyperbolic hexagon $\{ijk\}\in F$ on $\Omega_{ijk}(\eta)$, if $\lim f_{i}=-\infty$, then $\lim f_{j}=\lim f_{k}=+\infty$. Furthermore, $\lim\theta^{jk}_{i}(f_{i},f_{j},f_{k})=+\infty$.
@@
The definition of the admissible space $\mathcal{U}_{ijk}(\eta)$ in (35) implies $u_{r}+u_{s}>C(\eta_{rs})$ for any $\{r,s\}\subseteq\{i,j,k\}$. Hence, there exists no subset $\{r,s\}\subseteq\{i,j,k\}$ such that $\lim u_{r}=-\infty$ and either $\lim u_{s}=-\infty$ or $\lim u_{s}=c$, where $c$ is a constant. By combining (34) and Remark 2.22, the following hold: if $\lim f_{i}=-\infty$, then $\lim u_{i}=-\infty$; if $\lim f_{j}=+\infty$, then $\lim u_{j}=+\infty$ when $\alpha_{j}=0$ and $\lim u_{j}=0^{-}$ when $\alpha_{j}=1$. Therefore, if $\lim f_{i}=-\infty$, then $\lim u_{i}=-\infty$, which implies $\lim u_{j}=\lim u_{k}=+\infty$. This further yields $\alpha_{j}=\alpha_{k}=0$ and $\lim f_{j}=\lim f_{k}=+\infty$.
If $\alpha_{i}=0$, then $\cosh l_{ij}=-1+\eta_{ij}e^{u_{i}+u_{j}}$, $\cosh l_{jk}=-1+\eta_{jk}e^{u_{j}+u_{k}}$, and $\cosh l_{ik}=-1+\eta_{ik}e^{u_{i}+u_{k}}$. Thus
@@
Since $L_{2}>0$, it follows that $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, which implies $\theta^{jk}_{i}\rightarrow+\infty$.
@@
Similarly, $\cosh l_{ik}\rightarrow-1+2\eta_{ik}e^{u_{i}+u_{k}}$ and $\cosh l_{jk}=-1+\eta_{jk}e^{u_{j}+u_{k}}$. By reasoning analogous to the case $\alpha_{i}=0$, we have $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, which implies $\theta^{jk}_{i}\rightarrow+\infty$.
@@
Similarly, $\cosh l_{ik}\rightarrow-1+2\eta_{ik}e^{u_{i}+u_{k}}$ and $\cosh l_{jk}=-1+\eta_{jk}e^{u_{j}+u_{k}}$. Similar to the proof of the case $\alpha_{i}=0$, we have $\cosh\theta^{jk}_{i}=L_{1}+L_{2}\rightarrow+\infty$, which implies $\theta^{jk}_{i}\rightarrow+\infty$. Q.E.D.
Proof of Theorem 3.5: Let $X$ be the image of $K$. To prove $X=\mathbb{R}^{N}_{>0}$, we aim to show that $X$ is both an open and a closed subset of >$\mathbb{R}^{N}_{>0}$. By the definition of the curvature $K$, it follows directly that $X\subseteq\mathbb{R}^{N}_{>0}$. Since the map from $f\in\Omega(\eta)$ to >$\mathbb{R}^{N}_{>0}$ is a smooth injective map, its image $X$ must be an open subset of >$\mathbb{R}^{N}_{>0}$. It remains to verify that $X$ is closed in >$\mathbb{R}^{N}_{>0}$. Suppose $\{f^{(m)}\}$ is a sequence in $\Omega(\eta)$ such that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$. We shall show that there exists a subsequence, still denoted by $\{f^{(m)}\}$, such that $\lim_{m\rightarrow+\infty}f^{(m)}\in\Omega(\eta)$.
Suppose, for contradiction, that there exists a subsequence $\{f^{(m)}\}$ such that $\lim_{m\rightarrow+\infty}f^{(m)}$ lies on the boundary of $\Omega(\eta)$. By (33), the boundary of the admissible space $\Omega(\eta)$ in $[-\infty,+\infty]^{N}$ consists of three distinct components, specified as follows:
@@
There exists an ideal edge $\{rs\}\in E$ such that $l_{rs}=0$, a condition equivalent to $u_{r}+u_{s}=C(\eta_{rs})$ for the admissible space $\mathcal{U}_{ijk}(\eta)$ defined by (35);
@@
For $\alpha\in\{0,1\}$, there exists $i\in B$ such that $f_{i}^{(m)}\rightarrow+\infty$;
@@
For $\alpha\in\{0,1\}$, there exists $i\in B$ such that $f_{i}^{(m)}\rightarrow-\infty$.
For the case (i), consider a right-angled hyperbolic hexagon $\{ijk\}\in F$ with edge lengths $l_{ij},l_{jk},l_{ik}$ and opposite boundary lengths $\theta^{ij}_{k},\theta^{jk}_{i},\theta^{ik}_{j}$. Without loss of generality, we assume $l_{ij}\rightarrow 0$. By the cosine law, we have
@@
which contradicts the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$. For the case (ii), by Lemma 3.6, each generalized angle incident to the boundary component $i\in B$ converges to 0. Hence, $\lim_{m\rightarrow+\infty}K_{i}^{(m)}=0$. For the case (iii), Lemma 3.7 implies that $\lim_{m\rightarrow+\infty}K_{i}^{(m)}=+\infty$. Both cases contradict the assumption that $\lim_{m\rightarrow+\infty}K^{(m)}\in\mathbb{R}^{N}_{>0}$. Q.E.D.
@@
Note that Lemma 3.6 and Lemma 3.7 are valid for $\alpha:B\rightarrow\{-1,0,1\}$. In contrast, Theorem 3.5 only holds for $\alpha:B\rightarrow\{0,1\}$. Specifically, when $\alpha_{i}=-1$, the boundary of the admissible space $\Omega(\eta)$ in $[-\infty,+\infty]^{N}$ will include the case that there exists $i\in B$ such that $f_{i}^{(m)}\rightarrow 0^{-}$. For this case, we are unable to prove the closedness of $X$.